OpenStudy (anonymous):

Find an equation of the tangent line at the indicated point on the graph of the function. y = f(x) = x2 - x, (x, y) = (2, 2)

OpenStudy (anonymous):

To know the equation of a tangent line, you need two things. 1) You need the slope of the line. That's found by finding the derivative of the function and evaluating it at the point. 2) You also need a point that the tangent line goes through. But you already know that the tangent line passes through the point you're given. So you have that information as well. Knowing the point and slope, you can use the point-slope form of a line to write the equation of the tangent line.

OpenStudy (anonymous):

dy/dx=2x-1, it is the slope of line at any point, so at {2,2} it will be 3. now equation of line... (y-2)=3(x-2) 3x-y-4=0